CC12:

Triangle ABC has the following coordinates: A (-2, 3) B (4, -3) C (-2, -3) What are the coordinates points for triangle A'B'C' if the scale factor is −3/4.

Vocaloid:

You just have to multiple every coordinate by the scale factor -3/4

CC12:

-2,3*-3/4=-2,2.25 @dude

CC12:

\[4,-3\times-3/4=4,2.25\]

CC12:

\[-2,-3\times-3/4=-2,2.25\]

dude:

Well uh you want to watch your signs first negative times a negative is a positive \(\color{#0cbb34}{\text{Originally Posted by}}\) @CC12 -2,3*-3/4=-2,2.25 \(\color{#0cbb34}{\text{End of Quote}}\) -2\(\times-\frac34\) is 1.5 3\(\times-\frac34\) is -2.25 You're supposed to multiply the x and y value by the scale factor

CC12:

Oh i get it now

CC12:

do you add the 1.5 and 2.225 then

CC12:

2.25

dude:

No, its a coordinate so it stays as (1.5, 2.25)

dude:

That is point A Could you re-do the rest? (I'll check them)

CC12:

yes

CC12:

\[4\times-3/4=-3\]

CC12:

\[3\times-3/4=2.25\]

dude:

Right answer but when you show your work, make sure it corresponds to the numbers \(\color{#0cbb34}{\text{Originally Posted by}}\) @CC12 \[3\times-3/4=2.25\] \(\color{#0cbb34}{\text{End of Quote}}\) The 3 should be negative here, else it would give you -2.25

CC12:

-3*-3/4?

dude:

Ya

dude:

The answer was right

CC12:

-2*-3/4=1.5 -3*-3/4=2.25

dude:

Yep

CC12:

A'(1.5,2.25) B'(-3,2.25) C'(1.5,2.25)

CC12:

Thank you so much dude

dude:

Sure :)